a manufacturer estimates that D(p) = 3000e^-0.05p units of a particular good will be sold at market price of p cedis per unit. determine the market price that will result in marginal revenue of zero


1
Expert's answer
2022-02-02T12:34:40-0500

Market Price with Zero MR

Marginal Revenue is given as the change in total revenue divided by the change in total quantity.

Total revenue= "(D(p)\times P)"

change in total revenue "TR'(p)" on the other hand is given as;

"TR'(p)=\frac{d}{dp}(3000e^{-0.05p})=\frac{150}{-e^{z/20}}"

Marginal revenue "(MR=0)"

equating to zero we get;

"MR=-150e^{-0.05p}P^2 +3000e^{-0.05p}=0"

substituting,

"e^{-0.05p}\times(20-P^2)=0"

"e^{1\/20p}=0"

"P^2=20"

"P=4.47"

therefore, market price at zero MR is given as P=4.47